# SurgiATM: A Physics-Guided Plug-and-Play Model for Deep Learning-Based Smoke Removal in Laparoscopic Surgery

### v1

@@ Abstract
During laparoscopic surgery, smoke generated by tissue cauterization can significantly degrade the visual quality of endoscopic frames, increasing the risk of surgical errors and hindering both clinical decision-making and computer-assisted visual analysis. Consequently, removing surgical smoke is critical to ensuring patient safety and maintaining operative efficiency. In this study, we propose the Surgical Atmospheric Model (SurgiATM) for surgical smoke removal. SurgiATM statistically bridges a physics-based atmospheric model and data-driven deep learning models, combining the superior generalizability of the former with the high accuracy of the latter. Furthermore, SurgiATM is designed as a lightweight, plug-and-play module that can be seamlessly integrated into diverse surgical desmoking architectures to enhance their accuracy and stability, better meeting clinical requirements. It introduces only two hyperparameters and no additional trainable weights, preserving the original network architecture with minimal computational and modification overhead. We conduct extensive experiments on three public surgical datasets with ten desmoking methods, involving multiple network architectures and covering diverse procedures, including cholecystectomy, partial nephrectomy, and diaphragm dissection. The results demonstrate that incorporating SurgiATM commonly reduces the restoration errors of existing models and relatively enhances their generalizability, without adding any trainable layers or weights. This highlights the convenience, low cost, effectiveness, and generalizability of the proposed method. The code for SurgiATM is released at https://github.com/MingyuShengSMY/SurgiATM.
{IEEEkeywords}
@@ 1 Introduction
Minimally invasive surgery (MIS) has noticeably benefited patients by reducing risk and enabling faster recovery [Maier_Hein_2017_Surgical_data, Maier_Hein_2022_Surgical_data]. However, surgical smoke, commonly generated by high-energy instruments (e.g., electrosurgical units, laser surgical tools, and high-frequency ultrasonic scalpels), can severely impair endoscopic visibility, posing a notable challenge. The presence of dense smoke obstructs the surgeon’s field of view, increases the risk of surgical errors, and reduces procedural efficiency. Moreover, smoke-induced image degradation adversely affects vision-based downstream surgical tasks, such as phase recognition [Yang_2024_Surgformer_Surgical, Zou_2025_Capturing_action, Wagner_2023_Comparative_validation, Yue_2023_Cascade_Multi, Guan_2024_Label_Guided, Zhang_2026_CSAP_Assist], tool detection, semantic segmentation [Sheng_2025_AMNCutter_Affinity, Sheng_2024_Revisiting_Surgical, Yue_2024_SurgicalSAM_Efficient], depth estimation [Zou_2025_SSIFNet_Spatialtemporal, Tao_2023_SVT_SDE, Cheng_2022_Deep_Laparoscopic], 3D surface reconstruction [Gong_2024_Self_Supervised, Maier_Hein_2014_Comparative_Validation], etc., which require both high frame quality and real-time processing speed [Han_2025_Towards_Reliable, Long_2025_Surgical_embodied, Wang_2025_Learning_dissection, Xu_2026_Surgical_Action, Dou_2025_Artificial_intelligence] . At present, mechanical smoke evacuation or filtration is commonly achieved through hardware-based systems, which often require additional effort and impose extra workload during surgery, disrupting the surgical workflow and causing surgeon fatigue [Manning_2018_Laparoscopic_lens, Xia_2025_In_Vivo]. Therefore, ensuring a clear visual field in real time is critical [Ulmer_2008_The_hazards, Carbajo_Rodrguez_2009_Surgical_smoke, Weld_2007_Analysis_of].
@@
To this end, real-time image processing has gained increasing attention as a means of digitally removing smoke during intraoperative visualization in a more streamlined and automated way by leveraging computer vision and deep learning techniques. It offers a cost-effective, low-risk, and efficient alternative for surgical smoke removal [Luo_2017_Vision_Based, Chen_2020_De_smokeGCN, Salazar_Colores_2020_Desmoking_Laparoscopy, Li_2024_Multi_frequency, Wang_2024_Desmoking_of]. However, traditional models often struggle with this task, as they were originally designed for outdoor hazy environments. Unlike natural hazy images, in vivo surgical scenes present unique challenges, including complex physics parameter estimation due to non-uniform distribution of smoke, spatially heterogeneous illumination caused by radial light attenuation, inaccurate color restoration on non-smoke objects (e.g., surgical instruments), severe information loss resulting from highly dense smoke obscuration, and temporal incoherence of smoke occurrence due to the intermittent use of energy devices [Modrzejewski_2020_Light_modelling, Azagra_2023_Endomapper_dataset, Rodrguez_Puigvert_2023_LightDepth_Single, Stelzer_2024_Generation_and, Kumar_2023_Understanding_surgical, Tian_2015_Single_Image]. To address these limitations, deep learning-based approaches have been increasingly adopted for their powerful non-linear mapping capabilities and ease of deployment in laparoscopic systems [Chen_2022_DesmokeNet_A, Li_2022_Endoscopy_image]. These methods show considerable promise for surgical desmoking, offering more consistent color restoration and improved visual clarity. Nevertheless, due to the scarcity of paired desmoking benchmarks, most deep learning-based models face the challenge of generalizability, as they can produce high-quality predictions on a dataset with synthetic smoke while suffering from performance degradation on another dataset with real smoke. This is an important concern for real-world clinical applications [Wu_2025_Self_Supervised].
@@
  1. 1.
    We propose SurgiATM, a parameter-efficient, plug-and-play, physics-guided model, that improves performance and generalizability of surgical desmoking approaches to better meet clinical requirements.
  2. 2.
    The proposed model introduces zero trainable weights and can be seamlessly integrated into state-of-the-art (SOTA) methods built upon various backbone architectures (e.g., U-Net, Swin Transformer).
  3. 3.
    We validate SurgiATM on three public surgical datasets, demonstrating improvements in both performance and generalizability across multiple SOTA desmoking methods and various surgical scenes, thereby confirming its effectiveness, stability, and generalizability.
@@ 2.1 Physics-Based Methods
The Atmospheric Scattering Model (ASM) is a widely used physical model that describes the formation of hazy natural images by modeling the absorption, scattering, and reflection of light by objects and particles in the atmospheric medium [Narasimhan_2002_Vision_and, Narasimhan_2003_Contrast_restoration]. Based on the ASM, clear images can be restored from degraded observations by estimating the required physical parameters, such as medium transmission, atmospheric light, and scene depth, which has motivated extensive research in the field. In early studies, the Gaussian distribution and Markov Random Field (MRF) model were classic and well-known techniques for estimating the parameters [Fattal_2008_Single_image, Tan_2008_Visibility_in]. Then, [Kaiming_2011_Single_Image] proposed the Dark Channel Prior (DCP) model for natural image dehazing, in which the medium transmission and global atmospheric light were statistically estimated according to their physical properties. The DCP assumes that haze in an image is uniformly distributed and that, in hazeless images, at least one color channel in most non-sky patches or pixels has very low intensity. In contrast, haze tends to raise pixel intensities across all channels, causing the dark channel to become brighter and the image to appear more whitish. Since 2012, the field of surgical desmoking has gained increasing attention, with many valuable studies published over the past decade. [Gu_2015_Virtual_Digital] developed a digital defogging system that uses the DCP to enhance image quality in laparoscopic surgery. [Kotwal_2016_Joint_desmoking] combined the desmoking and denoising tasks of laparoscopic images and transformed them into a Bayesian inference problem by proposing a novel probabilistic graphical model. [Baid_2017_Joint_desmoking] extends [Kotwal_2016_Joint_desmoking] by introducing additional image priors of color and texture derived from sparse dictionary models. In [Tchaka_2017_Chromaticity_based], histogram equalization is combined with an adaptive DCP to remove surgical smoke. [Luo_2017_Vision_Based] rearranged the original atmospheric model by introducing an atmospheric veil to represent the transmission map and then solved the parameter estimation problem via the Poisson equation in the frequency domain. [Wang_2018_Variational_based] first developed an energy function to optimize and estimate a smoke veil to recover smokeless images based on [Tian_2015_Single_Image]. Combining the hardware camera and the software algorithm, [Azam_2022_Smoke_removal] proposed a method based on multiple-exposure fusion, which enhances local contrast and color saturation, thereby improving image quality. These works offer valuable insights into the feasibility of physics-based approaches for surgical desmoking, encouraging further research.
@@ 2.2 Deep Learning-Based Methods
To achieve higher restoration quality, the powerful fitting and generative capabilities of Deep Neural Networks (DNNs) offer an alternative solution [Cai_2016_DehazeNet_An, Li_2017_AOD_Net, Dong_2020_Multi_Scale, Qin_2020_FFA_Net, Bolkar_2018_Deep_Smoke, Sidorov_2020_Generative_Smoke]. Current deep learning-based desmoking methods can be broadly categorized into two frameworks: non-adversarial learning and adversarial learning.
@@ 2.2.1 Non-adversarial Learning Methods
This framework indicates that the model is tasked with learning a mapping from smoky or hazy images to their corresponding clean versions. [Chen_2018_Unsupervised_Learning] trained a U-Net model [Ronneberger_2015_U_Net] on laparoscopic videos with synthetic smoke using Blender [^1] https://www.blender.org/. [Bolkar_2018_Deep_Smoke] attempted to address surgical smoke removal via AODNet [Li_2017_AOD_Net], a Convolutional Neural Network (CNN) designed to estimate global light illuminance and the transmission map simultaneously. Similar to [Chen_2018_Unsupervised_Learning], [Bolkar_2018_Deep_Smoke] trained the model on synthetically generated smoky in vivo frames by blending Perlin noise with clean frames. Furthermore, instead of training on synthetic smoky frames, [Ma_2021_A_Smoke] attempted to use adjacent frames, along with the presence or absence of smoke, as training inputs and corresponding ground truth. Similarly, in the most recent study [Wu_2025_Self_Supervised], they implemented a real-world self-supervised video desmoking approach by capturing the informative features from frames recorded prior to the activation of high-energy surgical instruments. Based on the above training data strategy, considerable efforts have been made to improve performance further. [Wang_2019_Multiscale_deep] designed a U-Net-like network and proposed Laplacian image pyramid decomposition to capture multi-scale features. [Chen_2020_De_smokeGCN] presented a novel generative-collaborative learning framework (De-smokeGCN), dividing the desmoking into two sub-tasks: smoke detection and smoke removal. A smoke detector predicted the smoke mask, which was leveraged as a prior for the smoke removal network and motivated several subsequent studies [Hong_2023_MARS_GAN, Zhou_2024_Synchronizing_Detection]. [Lin_2021_A_desmoking] improved the decoder part of the U-Net by appending a convolutional block attention module to generate a guidance mask for feature maps. [Kanakatte_2021_Surgical_Smoke] introduced a workflow to simultaneously remove smoke and restore colors, using U-Net as the backbone. [Ma_2024_Laparoscopic_Video] leveraged deformable convolution incorporated with a mutual attention mechanism to model temporal features and combined Local Binary Patterns with input frames as a texture prior. [Wang_2024_Smoke_veil] developed a method consisting of three U-Nets to produce distinct outputs, each serving as an element in the energy function derived from their previous study [Wang_2018_Variational_based]. [Li_2024_Multi_frequency] developed a surgical desmoking method based on the Diffusion Model [Ho_2020_Denoising_Diffusion, Jiaming_2021_Denoising_Diffusion], incorporating a multi-level frequency analysis module to integrate features across different frequency bands. In addition to convolutional network architectures, the Swin Transformer (SwinT) [Liu_2021_Swin_Transformer] has recently gained increasing attention in this field because of its hierarchical representation and efficient local-to-global attention mechanism. [Wang_2023_Surgical_smoke] proposed a network that utilized several convolutional layers to extract low-level features and then used SwinT to analyze deep and global information from smoky frames. [Wang_2024_Desmoking_of] developed a U-shaped transformer model based on U-Net and SwinT to enhance feature representation by combining local detail extraction with global context modeling.
@@ 2.2.2 Adversarial Learning Methods
Apart from conventional image-to-image translation, converting laparoscopic images from the smoky domain to the smokeless domain offers an alternative paradigm for addressing the surgical desmoking problem. By regarding desmoking as an image style transfer task, the Generative Adversarial Network (GAN) [Goodfellow_2020_Generative_adversarial] and its variants have emerged as effective solutions. [Sidorov_2020_Generative_Smoke] modeled a mapping between smoky and smoke-free domains using a conventional GAN as the backbone and leveraging the perceptual image quality score as a loss function to improve the result quality. [Salazar_Colores_2020_Desmoking_Laparoscopy] developed a method based on a Conditional Generative Adversarial Network (cGAN) [Mirza_2014_Conditional_generative] and used an embedded dark channel as a prior. Compared to the above GANs, the Cycle-Consistent Adversarial Network (CycleGAN) [Zhu_2017_Unpaired_Image] is well-suited for image style transfer tasks and is therefore more commonly used for surgical smoke removal [Su_2023_Multi_stages, Wang_2025_Desmoke_VCU, Islam_2024_LVQE_Laparoscopic]. [Vishal_2019_Guided_Unsupervised] enhanced CycleGAN by incorporating multi-scale feature extraction and introducing an upsampling loss to improve the contrast of the desmoked frames. In their subsequent studies, they further included structure-consistency loss and designed a refinement module to improve restoration quality [Vishal_2020_Unsupervised_Desmoking, Venkatesh_2020_Unsupervised_smoke]. [Pan_2022_DeSmoke_LAP] further designed internal-channel and dark-channel loss functions based on the characteristics of smoky pixels to improve model performance. Similar to [Chen_2020_De_smokeGCN], [Zhou_2024_Synchronizing_Detection] added an extra detection network to the CycleGAN architecture to estimate the smoke mask as a prior for subsequent desmoking. [Hong_2023_MARS_GAN] adopted a multilevel strategy to adaptively learn non-homogeneous smoke features, using the predicted smoke distribution as a prior.

### v2

@@ Abstract
During laparoscopic surgery, smoke generated by tissue cauterization can significantly degrade the visual quality of endoscopic frames, increasing the risk of surgical errors and hindering both clinical decision-making and computer-assisted visual analysis. Consequently, removing surgical smoke is critical to ensuring patient safety and maintaining operative efficiency. In this study, we propose the Surgical Atmospheric Model (SurgiATM) for surgical smoke removal. SurgiATM statistically bridges a physics-based atmospheric model and data-driven deep learning models, combining the superior generalizability of the former with the high accuracy of the latter. Furthermore, SurgiATM is designed as a lightweight, plug-and-play module that can be seamlessly integrated into diverse surgical desmoking architectures to enhance their accuracy and stability. The proposed method is derived via statistically optimizing Mixture-of-Experts (MoE) model at the output end of arbitrary deep learning methods, with a Laplacian-like error distribution specifically leveraged to model surgical smoke. The output-stage MoE ensures minimal modification to the architecture of the original methods, while the Laplacian-like distribution characteristic of surgical smoke enables a lightweight reconstruction formulation with minimal parameters. Therefore, SurgiATM introduces only two hyperparameters and no additional trainable weights, preserving the original network architecture with minimal computational and modification overhead. We conduct extensive experiments on three public surgical datasets with ten desmoking methods, involving multiple network architectures and covering diverse procedures, including cholecystectomy, partial nephrectomy, and diaphragm dissection. The results demonstrate that incorporating SurgiATM commonly reduces the restoration errors of existing models and relatively enhances their generalizability, without adding any trainable layers or weights. This highlights the convenience, low cost, effectiveness, and generalizability of the proposed method. The code for SurgiATM is released at https://github.com/MingyuShengSMY/SurgiATM.
keywords:
@@
journal: Medical Image Analysis
[…]
[usyd]organization=School of Computer Science, The University of Sydney,city=Sydney, state=NSW, country=Australia
[…]
[sjt]organization=Institute of Medical Robotics, School of Biomedical Engineering, Shanghai Jiao Tong University,city=Shanghai, country=China
[…]
[hvd]organization=Department of Radiology, Brigham and Women’s Hospital, and Harvard Medical School,city=Boston, state=MA, country=United States
@@ 1 Introduction
Minimally invasive surgery (MIS) has noticeably benefited patients by reducing risk and enabling faster recovery (Maier-Hein et al., 2017, 2022). However, surgical smoke, commonly generated by high-energy instruments (e.g., electrosurgical units, laser surgical tools, and high-frequency ultrasonic scalpels), can severely impair endoscopic visibility, posing a notable challenge (Duelmer et al., 2026). The presence of dense smoke obstructs the surgeon’s field of view, increases the risk of surgical errors, and reduces procedural efficiency. Moreover, smoke-induced image degradation adversely affects vision-based downstream surgical tasks, such as phase recognition (Yang et al., 2024; Zou et al., 2025a; Wagner et al., 2023; Yue et al., 2023; Guan et al., 2024; Zhang et al., 2025; Das et al., 2025), tool detection, semantic segmentation (Sheng et al., 2025, 2024; Yue et al., 2024; Sestini et al., 2025; Cartucho et al., 2024), depth estimation (Zou et al., 2025b; Tao et al., 2023; Cheng et al., 2022), 3D surface reconstruction (Gong et al., 2024; Maier-Hein et al., 2014; Rau et al., 2024), etc., which require both high frame quality and real-time processing speed (Han et al., 2025; Long et al., 2025; Wang et al., 2025a; Xu et al., 2026; Dou et al., 2025; Rueckert et al., 2026). At present, mechanical smoke evacuation or filtration is commonly achieved through hardware-based systems, which often require additional effort and impose extra workload during surgery, disrupting the surgical workflow and causing surgeon fatigue (Manning et al., 2018; Xia et al., 2025). Therefore, ensuring a clear visual field in real time is critical (Ulmer, 2008; Carbajo-Rodríguez et al., 2009; Weld et al., 2007).
@@
To this end, real-time image processing has gained increasing attention as a means of digitally removing smoke during intraoperative visualization in a more streamlined and automated way by leveraging computer vision and deep learning techniques. It offers a cost-effective, low-risk, and efficient alternative for surgical smoke removal (Luo et al., 2017; Chen et al., 2020; Salazar-Colores et al., 2020; Li et al., 2024; Wang et al., 2024b). However, traditional models often struggle with this task, as they were originally designed for outdoor hazy environments. Unlike natural hazy images, in vivo surgical scenes present unique challenges, including complex physics parameter estimation due to non-uniform distribution of smoke, spatially heterogeneous illumination caused by radial light attenuation, inaccurate color restoration on non-smoke objects (e.g., surgical instruments), severe information loss resulting from highly dense smoke obscuration, and temporal incoherence of smoke occurrence due to the intermittent use of energy devices (Modrzejewski et al., 2020; Azagra et al., 2023; Rodríguez-Puigvert et al., 2023; Stelzer et al., 2024; Kumar et al., 2023; Tian et al., 2015; Ramesh et al., 2023). To address these limitations, deep learning-based approaches have been increasingly adopted for their powerful non-linear mapping capabilities and ease of deployment in laparoscopic systems (Chen et al., 2022; Li et al., 2022). These methods show considerable promise for surgical desmoking, offering more consistent color restoration and improved visual clarity. Nevertheless, due to the scarcity of paired desmoking benchmarks, most deep learning-based models face the challenge of generalizability, as they can produce high-quality predictions on a dataset with synthetic smoke while suffering from performance degradation on another dataset with real smoke. This is an important concern for real-world clinical applications (Wu et al., 2025b).
@@
  • We propose SurgiATM, a parameter-efficient, plug-and-play, physics-guided model, that improves performance and generalizability of surgical desmoking approaches, providing clearer desmoking results.
  • The proposed model is derived via Mixture-of-Experts (MoE) model at the output end, and thus SurgiATM can be employed as an output layer (or module), seamlessly integrated into state-of-the-art (SOTA) methods built upon various backbone architectures (e.g., U-Net, Swin Transformer).
  • We develop SurgiATM based on statistical analysis and physical model; thereby it introduces zero trainable weights, while only two hyper-parameters are involved. The recommended configurations for the two hyper-parameters are analyzed in ablation study.
  • We validate SurgiATM on three public surgical datasets, demonstrating improvements in both performance and generalizability across multiple SOTA desmoking methods and various laparoscopic surgical scenes, thereby confirming its effectiveness, low efficiency overhead, stability, and generalizability for laparoscopic surgical smoke removal.
@@ 2.1 Physics-Based Methods
The Atmospheric Scattering Model (ASM) is a widely used physical model that describes the formation of hazy natural images by modeling the absorption, scattering, and reflection of light by objects and particles in the atmospheric medium (Narasimhan and Nayar, 2002, 2003). Based on the ASM, clear images can be restored from degraded observations by estimating the required physical parameters, such as medium transmission, atmospheric light, and scene depth, which has motivated extensive research in the field. In early studies, the Gaussian distribution and Markov Random Field (MRF) model were classic and well-known techniques for estimating the parameters (Fattal, 2008; Tan, 2008). Then, Kaiming et al. (2011) proposed the Dark Channel Prior (DCP) model for natural image dehazing, in which the medium transmission and global atmospheric light were statistically estimated according to their physical properties. The DCP assumes that haze in an image is uniformly distributed and that, in hazeless images, at least one color channel in most non-sky patches or pixels has very low intensity. In contrast, haze tends to raise pixel intensities across all channels, causing the dark channel to become brighter and the image to appear more whitish. Since 2012, the field of surgical desmoking has gained increasing attention, with many valuable studies published over the past decade. Gu et al. (2015) developed a digital defogging system that uses the DCP to enhance image quality in laparoscopic surgery. Kotwal et al. (2016) combined the desmoking and denoising tasks of laparoscopic images and transformed them into a Bayesian inference problem by proposing a novel probabilistic graphical model. Baid et al. (2017) extends (Kotwal et al., 2016) by introducing additional image priors of color and texture derived from sparse dictionary models. In Tchaka et al. (2017) study, histogram equalization is combined with an adaptive DCP to remove surgical smoke. Luo et al. (2017) rearranged the original atmospheric model by introducing an atmospheric veil to represent the transmission map and then solved the parameter estimation problem via the Poisson equation in the frequency domain. Wang et al. (2018) first developed an energy function to optimize and estimate a smoke veil to recover smokeless images based on Tian et al. (2015). Combining the hardware camera and the software algorithm, Azam et al. (2022) proposed a method based on multiple-exposure fusion, which enhances local contrast and color saturation, thereby improving image quality. These works offer valuable insights into the feasibility of physics-based approaches for surgical desmoking, encouraging further research.
@@ 2.2 Deep Learning-Based Methods
To achieve higher restoration quality, the powerful fitting and generative capabilities of Deep Neural Networks (DNNs) offer an alternative solution (Cai et al., 2016; Li et al., 2017; Dong et al., 2020; Qin et al., 2020; Bolkar et al., 2018; Sidorov et al., 2020). Current deep learning-based desmoking methods can be broadly categorized into two frameworks: non-adversarial learning and adversarial learning.
@@ 2.2.1 Non-adversarial Learning Methods
This framework indicates that the model is tasked with learning a mapping from smoky or hazy images to their corresponding clean versions. Chen et al. (2018) trained a U-Net model (Ronneberger et al., 2015) on laparoscopic videos with synthetic smoke using Blender [^1] https://www.blender.org/. Bolkar et al. (2018) attempted to address surgical smoke removal via AODNet (Li et al., 2017), a Convolutional Neural Network (CNN) designed to estimate global light illuminance and the transmission map simultaneously. Similar to Chen et al. (2018), Bolkar et al. (2018) trained the model on synthetically generated smoky in vivo frames by blending Perlin noise with clean frames. Furthermore, instead of training on synthetic smoky frames, Ma et al. (2021) attempted to use adjacent frames, along with the presence or absence of smoke, as training inputs and corresponding ground truth. Similarly, in the most recent study, Wu et al. (2025b) implemented a real-world self-supervised video desmoking approach by capturing the informative features from frames recorded prior to the activation of high-energy surgical instruments. Based on the above training data strategy, considerable efforts have been made to improve performance further. Wang et al. (2019) designed a U-Net-like network and proposed Laplacian image pyramid decomposition to capture multi-scale features. Chen et al. (2020) presented a novel generative-collaborative learning framework (De-smokeGCN), dividing the desmoking into two sub-tasks: smoke detection and smoke removal. A smoke detector predicted the smoke mask, which was leveraged as a prior for the smoke removal network and motivated several subsequent studies (Hong et al., 2023; Zhou et al., 2024). Lin et al. (2021) improved the decoder part of the U-Net by appending a convolutional block attention module to generate a guidance mask for feature maps. Kanakatte et al. (2021) introduced a workflow to simultaneously remove smoke and restore colors, using U-Net as the backbone. Ma et al. (2024a) leveraged deformable convolution incorporated with a mutual attention mechanism to model temporal features and combined Local Binary Patterns with input frames as a texture prior. Wang et al. (2024a) developed a method consisting of three U-Nets to produce distinct outputs, each serving as an element in the energy function derived from their previous study (Wang et al., 2018). Li et al. (2024) developed a surgical desmoking method based on the Diffusion Model (Ho et al., 2020; Jiaming et al., 2021), incorporating a multi-level frequency analysis module to integrate features across different frequency bands. Wu et al. (2025a) further combine the information in frequency and temporal domains, and developed a multi-scale Gaussian window constraint to capture inter-frame features with fine structural details preserved. In addition to convolutional network architectures, the Swin Transformer (SwinT) (Liu et al., 2021) has recently gained increasing attention in this field because of its hierarchical representation and efficient local-to-global attention mechanism. Wang et al. (2023) proposed a network that utilized several convolutional layers to extract low-level features and then used SwinT to analyze deep and global information from smoky frames. Wang et al. (2024b) developed a U-shaped transformer model based on U-Net and SwinT to enhance feature representation by combining local detail extraction with global context modeling.
Figure 1: The proposed SurgiATM is derived from Atmospheric Scattering Model, Normalized Radiance Decomposition, Mixture-of-Experts (MoE), and Laplacian Error Minimization. We collect prediction results and errors from a lot of SOTA methods with diverse backbone architectures; and then Laplacian error distribution is observed from surgical desmoking which is different from natural fog. Compared to physics-based methods (e.g., DCP), data-driven methods usually demonstrate higher accuracy, while the formers show strong stability. To complement their strengths, SurgiATM is developed by optimizing this system.
@@ 2.2.2 Adversarial Learning Methods
Apart from conventional image-to-image translation, converting laparoscopic images from the smoky domain to the smokeless domain offers an alternative paradigm for addressing the surgical desmoking problem. By regarding desmoking as an image style transfer task, the Generative Adversarial Network (GAN) (Goodfellow et al., 2020) and its variants have emerged as effective solutions. Sidorov et al. (2020) modeled a mapping between smoky and smoke-free domains using a conventional GAN as the backbone and leveraging the perceptual image quality score as a loss function to improve the result quality. Salazar-Colores et al. (2020) developed a method based on a Conditional Generative Adversarial Network (cGAN) (Mirza and Osindero, 2014) and used an embedded dark channel as a prior. Compared to the above GANs, the Cycle-Consistent Adversarial Network (CycleGAN) (Zhu et al., 2017) is well-suited for image style transfer tasks and is therefore more commonly used for surgical smoke removal (Su and Wu, 2023; Wang et al., 2025b; Islam et al., 2024). Vishal et al. (2019) enhanced CycleGAN by incorporating multi-scale feature extraction and introducing an upsampling loss to improve the contrast of the desmoked frames. In their subsequent studies, they further included structure-consistency loss and designed a refinement module to improve restoration quality (Vishal et al., 2020; Venkatesh et al., 2020). Pan et al. (2022) further designed internal-channel and dark-channel loss functions based on the characteristics of smoky pixels to improve model performance. Similar to Chen et al. (2020), Zhou et al. (2024) added an extra detection network to the CycleGAN architecture to estimate the smoke mask as a prior for subsequent desmoking. Hong et al. (2023) adopted a multilevel strategy to adaptively learn non-homogeneous smoke features, using the predicted smoke distribution as a prior.
@@
Table 1: Primary Notations.
Variable Definition
II Smoky image
$J$ Smoke-free image
$_{\text{DCP}},J_{\text{DNN}} JJ$ predicted by DCP or DNN
$t$ Medium transmission
AA Global atmospheric light
$\rho$ Normalized radiance
$D$ Dark channel
$\mathcal{D}$ Denormalized dark channel
$W$ Expert-weights for MoE
$\varepsilon$ Error term
$\mu,b$ Laplacian parameters
$\eta$ Smoothing factor
$z$ Window size